OpenStudy (anonymous):

Given f(x) = 3x − 10 and g(x) = 12x + 8, solve for (f + g)(x) and select the correct answer below. (f + g)(x) = 15x − 2 (f + g)(x) = 9x − 2 (f + g)(x) = 15x + 2 (f + g)(x) = 36x2 − 120x

OpenStudy (anonymous):

cAN SOMEONE HELP??

OpenStudy (anonymous):

i got 47\11 before but its not even an answer choice

OpenStudy (nincompoop):

add f(x) and g(x)

OpenStudy (nincompoop):

f(x) = 3x − 10; g(x) = 12x + 8 (f + g) (x)) = 3x -10 +12x + 8

OpenStudy (anonymous):

(f+g)(x)= 15x-2 ??

OpenStudy (nincompoop):

:) that is all you have to do, combine them using addition

OpenStudy (anonymous):

yea i tried before, and failed, i got 47 over 11 haha

OpenStudy (nincompoop):

I do not know what you did before, but you seem to get it now

OpenStudy (anonymous):

Thankyou :)